OpenStudy (anonymous):

You are tasked with making a microscope with a magnification of 1000 using a converging lens with a focal distance of 1.5 cm for the objective. The tube length given to you is 30 cm long. What should the focal distance of the eyepiece be if you want the eyes to be relaxed while observing with this microscope? Assume a near point of 25 cm.

OpenStudy (anonymous):

For compound microscope: If you want the eye to be relaxed then magnification of the miocroscope=(Vo/Uo)(D/fe) i.e (Vo/Uo)(D/fe)=1000.....(1) and for objective lens : apply lens formula 1/Vo-1/Uo=i/fo 1-Vo/U0=Vo/fo Vo/Uo=1-Vo/fo...(2) from (1) and (2) we get.... (1-Vo/fo)(D/fe)=1000....(3) since length of the tube is 30cm then Vo+fe=30 cm....(4) from (3)and (4), get the value of fe......